OpenStudy (anonymous):

Express the complex number in trigonometric form. -2i

OpenStudy (freckles):

-2i=-2(i)=-2(0+i) =-2(0+i*1) so when is cos 0 and sin 1 ? or look at this way -2i=2(-i)=2(0-i) =2(0-1*i) so when is cos 0 and sin -1?

OpenStudy (freckles):

yes so for the bottom route you found theta to be 3pi/2

OpenStudy (anonymous):

pi/2 or 3pi/2 ?

OpenStudy (freckles):

\[2(\cos(\frac{3 \pi}{2})+i \sin(\frac{3\pi}{2}))\]

OpenStudy (anonymous):

Ok thank you

OpenStudy (freckles):

you could have went the top route too \[-2(\cos(\frac{\pi}{2})+i \sin(\frac{\pi}{2}))\]

OpenStudy (anonymous):

Can you help with one more? Express the complex number in trigonometric form. (2 points) 3 - 3i

OpenStudy (freckles):

well first both terms have a three in common

OpenStudy (freckles):

3(1-i)

OpenStudy (freckles):

when are sin and cos opposites?

OpenStudy (freckles):

like you have cos is pos and sin is neg so that means you are in the fourth quadrant what angle in the fourth quadrant gives you sin and cos are not only difference in sign but are opposite values

OpenStudy (anonymous):

2nd & 4th quadrants?

OpenStudy (anonymous):

oh I think I got it

OpenStudy (anonymous):

it would be 7pi/4 since we're in the 4th quadrant

OpenStudy (freckles):

great you see at 7pi/4 cos is -sqrt(2)/2 and sin is sqrt(2)/2 so but we have 3(1-i) we can manipulate that \[3 \frac{2}{ \sqrt{2}}( \frac{\sqrt{2}}{2}+i (-\frac{\sqrt{2}}{2})) \\ 3 \frac{2}{\sqrt{2}}(\cos(\frac{7 \pi}{4})+i \sin(\frac{7 \pi}{4}))\]

OpenStudy (freckles):

\[\frac{6}{\sqrt{2}}(\cos(\frac{7 \pi}{4})+i \sin(\frac{7 \pi}{4}))\]

OpenStudy (anonymous):

Would the answer be 3*sqr rt(2) (cos7pi/4 + isin7pi/4)

OpenStudy (freckles):

if you choose to rationalize the outside factor's denominator yep

OpenStudy (anonymous):

Ok also in the beginning how did you know that the cosine and sine would be opposites? Like in order to find out what quadrant it would be in.

OpenStudy (freckles):

we have 1-1i

OpenStudy (freckles):

1 and -1 are opposites

OpenStudy (freckles):

correcting type-o above: great you see at 7pi/4 cos is sqrt(2)/2 and sin is -sqrt(2)/2

OpenStudy (anonymous):

Oh okay

OpenStudy (freckles):

if you don't like that I could show you another way

OpenStudy (anonymous):

No I understand. Thank you

OpenStudy (freckles):

3-3i draw visual   \[3 \sqrt{2}(\cos(\frac{-\pi}{4})+i \sin(\frac{-\pi}{4}))\] this answer looks a little difference because I describe theta a little differently but the trig forms or still equal